A teacher allows the students to each "throw out" their lowest test scores in order to increase their overall grade. The original mean score of 5 test is 75.2. After the removal of the lowest test, the mean score becomes 80.25. What was the score of the removed test?
Answer by stanbon(75887)

A teacher allows the students to each "throw out" their lowest test scores in order to increase their overall grade. The original mean score of 5 test is 75.2. After the removal of the lowest test, the mean score becomes 80.25. What was the score of the removed test?
------------------------------------
Original Point total = 5*75.2 = 376
-----------
Let the removed score be "x".
----------------------------------
New mean score Equation: (376-x)/4 = 80.25
376-x = 321
x = 376-321
x = 55 (the low score that was removed)
=============================================
Cheers,
Stan H.